differentiate the following function with respect to x
(sin^-1(x))^x

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

Y=(sin-1x)x

taking log on both sides

logy=xlogsin-1x

(1/y )dy/dx=x.1/sin-1x.1/√(1-x2) +logsin-1x .1

dy/dx=(sin-1x)x[x/{sin-1x√(1-x2)}+logsin-1x]

AskIITians Faculty

Rinkoo Gupta

Thanks & Regards